if x + 3y =12 and 2/3 x -y =5 what is the value of x
thanks

Answer by mananth(16949)

if x + 3y =12 and 2/3 x -y =5 what is the value of x
x +3y=12
2/3 x -y=5
2x-3y /3= 5
2x-3y =15
x+3y+2x-3y= 12+15
3x=27
x=9